# Mean-field limits for Consensus-Based Optimization and Sampling

### v3

@@ Conclusion
A crucial detail here is the following: Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below. This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d. mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ does not need to be included in the definition of $\Omega_{J,t}$. This trick allows to show the strong mean-field limit result (2.12), improving on the previous results from [CBO-mfl-Huang2021, CBO-hypersurface-Fornasier_2020, fornasier2021consensusbased].
@@ 2.5 Quantitative mean-field limit for CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of sufficiently high order. Consider the systems (2.8) and (2.9) with the coefficients given by
$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.18)
@@
In the same way as for CBO, it is possible to prove the following more general statement. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$. If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
@@
    $\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
@@
    $\eta:=\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
@@ 3.1 Wasserstein stability estimates for the weighted moments
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant L_{\mathcal{M}}\,W_{p}(\mu,\nu).$ (3.1)
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L_{\mathcal{C}}\,W_{p}(\mu,\nu).$ (3.2)
@@
$\mathcal{M}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]},\qquad g(x):=\operatorname{e}^{-\beta f(x)}x,\qquad h(x):=\operatorname{e}^{-\beta f(x)}.$
@@
$\leqslant L_{\mathcal{M}}\,\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\,W_{p}(\mu,\nu)$
$\leqslant L_{\mathcal{M}}^{2}\,\Bigl{(}W_{p}(\mu,\delta_{0})+W_{p}(\nu,\delta_{0})\Bigr{)}\,W_{p}(\mu,\nu)\leqslant 2L_{\mathcal{M}}^{2}R\,W_{p}(\mu,\nu).$
@@
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\frac{\mu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\mu[\operatorname{e}^{-\beta f(x)}]}-\frac{\nu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\nu[\operatorname{e}^{-\beta f(x)}]}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu),$
@@
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant L_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$ (3.5)
@@
Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$. By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$. Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$. Then
@@
The first term is bounded from above by $CW_{p}(\mu^{J},\nu^{J})$ by the base case, while the other two terms converge to 0 in the limit as $J\to\infty$ by Proposition 3.1. Taking the limit $J\to\infty$, we deduce that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant CW_{p}(\mu,\nu).$
@@
$W_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\Bigr{\rVert}_{\rm F}\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,W_{p}(\mu,\nu).$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)-\mathcal{T}(\nu)\bigr{\rVert}\leqslant L_{\mathcal{T}}(R)\,W_{p}(\mu,\nu).$ (3.8)
@@
    $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant KW_{p}(\mu,\delta_{0}).$ (3.9)
@@
$\forall i\in\mathbf{N},\qquad\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\geqslant i.$ (3.11)
In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8). Therefore, $W_{p}(\nu_{i},\delta_{0})\to\infty$. On the other hand, by the triangle inequality,
$W_{p}\left(\mu_{i},\nu_{i}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-W_{p}\left(\mu_{i},\delta_{0}\right)\geqslant W_{p}\left(\nu_{i},\delta_{0}\right)-R.$
@@
$\limsup_{i\to\infty}\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{W_{p}\left(\mu_{i},\nu_{i}\right)}\leqslant K\limsup_{i\to\infty}\frac{W_{p}(\nu_{i},\delta_{0})+R}{W_{p}\left(\nu_{i},\delta_{0}\right)-R}=K,$

### v4

@@ Conclusion
A crucial detail here is that Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below. This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d. mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ does not need to be included in the definition of $\Omega_{J,t}$. This trick allows to show the strong mean-field limit result (2.12), improving on the previous results from [CBO-mfl-Huang2021, CBO-hypersurface-Fornasier_2020, fornasier2021consensusbased].
@@ 2.5 Quantitative mean-field limit for CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of all orders. Consider the systems (2.8) and (2.9) with the coefficients given by
$b(x,\mu)\coloneq-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu)\coloneq\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}(\mu)}.$ (2.18)
@@
In the same way as for CBO, it is possible to prove the following more general statement. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$. Assume that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$. If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$
@@
    $\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}-\sqrt{\mathcal{C}_{\beta}(\mu)}\right\rVert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}\coloneq\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
@@
    $\eta\coloneq\inf_{t\in[0,T]}\lambda_{\text{min}}\bigl{(}\mathcal{C}_{\beta}(\overline{\rho}_{t})\bigr{)}>0.$
@@
2.6 Numerical experiment
In this section, we present a numerical experiment illustrating our first main result, on the quantitative mean field limit for CBO; see Theorem 2.6. To this end, we apply CBO to minimize the Ackley function, a standard benchmark function in optimization given by
[…]
where $d$ is the dimension. This function, which is illustrated in Figure 2, is globally Lipschitz continuous and bounded from above and below uniformly over d, and so $f_{\rm Ack}\in\mathcal{A}(0,0,0)$. To illustrate the theorem, we first approximate the mean field dynamics by simulating the interacting particle system (1.1) using a large number $J_{\infty}$ of particles. Then, for various smaller values of $J\ll J_{\infty}$, we simulate the CBO evolution of a system of $J$ particles, initialized at the same positions and driven by the same Brownian motions as the first $J$ particles of the large system of size $J_{\infty}$. In all cases, we use CBO with isotropic noise, discretized using the Euler–Maruyama method with time step $\Delta t$. All the parameters used in this experiment are given in Table 1.
To exhibit the scaling as $J^{-\frac{1}{2}}$ in Theorem 2.6, we fix $p=2$ and approximate the left-hand side of (2.11) using a Monte Carlo approach with $M$ independent realizations. More precisely, omitting discretization for simplicity, we use the following estimator $\widehat{E}(J)$ of $E(J)$, where
$E(J)\coloneq\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2}\right],\qquad\widehat{E}(J)\coloneq\frac{1}{MJ}\sum_{m=1}^{M}\sum_{j=1}^{J}\sup_{t\in[0,T]}\Big{\lvert}X^{(j/J,m)}_{t}-X^{(j/J_{\infty},m)}_{t}\Big{\rvert}^{2}.$
Here $X^{(j/J,m)}_{t}$ denotes the position of the $j$-th particle in the system of size $J$, for the $m$-th independent realization of the particle systems.
Parameter Notation Value
$\int$ Dimension $d$ 2
$\int$ Inverse temperature $\beta$ $3$
$\int$ Diffusion coefficient $\sigma$ $\frac{1}{5}$
$\int$ Initial law $\overline{\rho}_{0}$ $\mathcal{N}(0,\operatorname{I}_{2})$
$\int$ Time step $\Delta t$ 0.01
$\int$ Final time $T$ 1
$\int$ Size for mean field approximation $J_{\infty}$ $10^{6}$
$\int$ Number of independent simulations $M$ $100$
Table 1: Parameters of the numerical experiment described in Subsection 2.6.
The results of the numerical experiment are illustrated in the right panel Figure 2. The expected scaling $E(J)\propto\frac{1}{J}$ is clearly observed for large values of $J$. We note, however, that this scaling of the error holds only for relatively large $J$. For small $J$, the effective number of particles contributing to the weighted average, as quantified by the effective sample size [MR3696003], is close to 1 in the early stages of the simulation. Therefore, the Monte Carlo scaling is not expected in this regime.
Figure 2: Left: Ackley function. Right: Monte Carlo estimator $\widehat{E}(J)$ of $E(J)$ for $J\in\{10\cdot 2^{k}\}_{k=0}^{14}$, and straight line exhibiting the asymptotic scaling as $J^{-1}$ of $E(J)$. Error bars corresponding to two standard deviations are depicted on both sides of the Monte Carlo estimator.
@@ 3.1 Wasserstein stability estimates for the weighted moments
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\rvert\leqslant L_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.1)
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\Bigl{\lVert}\mathcal{C}_{\beta}(\mu)-\mathcal{C}_{\beta}(\nu)\Bigr{\rVert}_{\rm F}\leqslant L_{\mathcal{C}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.2)
@@
$\mathcal{M}_{\beta}(\mu)=\frac{\mu[g]}{\mu[h]},\qquad g(x):=\operatorname{e}^{-\beta f(x)}x,\qquad h(x)\coloneq\operatorname{e}^{-\beta f(x)}.$
@@
$\leqslant L_{\mathcal{M}}\,\Bigl{(}\left\lvert\mathcal{M}_{\beta}(\mu)\right\rvert+\left\lvert\mathcal{M}_{\beta}(\nu)\right\rvert\Bigr{)}\,\mathcal{W}_{p}(\mu,\nu)$
$\leqslant L_{\mathcal{M}}^{2}\,\Bigl{(}\mathcal{W}_{p}(\mu,\delta_{0})+W_{p}(\nu,\delta_{0})\Bigr{)}\,\mathcal{W}_{p}(\mu,\nu)\leqslant 2L_{\mathcal{M}}^{2}R\,\mathcal{W}_{p}(\mu,\nu).$
@@
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\frac{\mu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\mu[\operatorname{e}^{-\beta f(x)}]}-\frac{\nu[\operatorname{e}^{-\beta f(x)}x\otimes x]}{\nu[\operatorname{e}^{-\beta f(x)}]}\right\rVert_{\rm F}\leqslant C\mathcal{W}_{p}(\mu,\nu),$
@@
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)},\qquad\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant L_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$ (3.5)
@@
Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$. By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $\mathcal{W}_{p}(\mu^{J},\mu)\to 0$ and $\mathcal{W}_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$. Since convergence in $\mathcal{W}_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$. Then
@@
The first term is bounded from above by $C\mathcal{W}_{p}(\mu^{J},\nu^{J})$ by the base case, while the other two terms converge to 0 in the limit as $J\to\infty$ by Proposition 3.1. Taking the limit $J\to\infty$, we deduce that
$\left\lVert\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\right\rVert_{\rm F}\leqslant C\mathcal{W}_{p}(\mu,\nu).$
@@
$\mathcal{W}_{p}(\mu^{J},\nu^{J})=\min_{\sigma\in\mathcal{S}_{J}}\left(\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}-Y^{\sigma(j)}\right\rvert^{p}\right)^{\frac{1}{p}},$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p}(^{d}),\qquad\Bigl{\lVert}\sqrt{\mathcal{C}_{\beta}(\mu)}-\sqrt{\mathcal{C}_{\beta}(\nu)}\Bigr{\rVert}_{\rm F}\leqslant\widetilde{L}_{\sqrt{\mathcal{C}}}\,\mathcal{W}_{p}(\mu,\nu).$
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)-\mathcal{T}(\nu)\bigr{\rVert}\leqslant L_{\mathcal{T}}(R)\,\mathcal{W}_{p}(\mu,\nu).$ (3.8)
@@
    $\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\bigl{\lVert}\mathcal{T}(\mu)\bigr{\rVert}\leqslant K\mathcal{W}_{p}(\mu,\delta_{0}).$ (3.9)
@@
$\forall i\in\mathbf{N},\qquad\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)}\geqslant i.$ (3.11)
In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of (3.11) would be bounded from above uniformly in $i$ by (3.8). Therefore, $\mathcal{W}_{p}(\nu_{i},\delta_{0})\to\infty$. On the other hand, by the triangle inequality,
$\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)\geqslant\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-\mathcal{W}_{p}\left(\mu_{i},\delta_{0}\right)\geqslant\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-R.$
@@
$\limsup_{i\to\infty}\frac{\Bigl{\lVert}\mathcal{T}(\mu_{i})-\mathcal{T}(\nu_{i})\Bigr{\rVert}}{\mathcal{W}_{p}\left(\mu_{i},\nu_{i}\right)}\leqslant K\limsup_{i\to\infty}\frac{\mathcal{W}_{p}(\nu_{i},\delta_{0})+R}{\mathcal{W}_{p}\left(\nu_{i},\delta_{0}\right)-R}=K,$